Weston Products manufactures an industrial cleaning compound that goes through three processing departments—Grinding, Mixing,Required 1 Required 2 Required 3 Required 4 Compute the Grinding Departments equivalent units of production for materials anRequired 1 Required 2 Required 3 Required 4 Compute the Grinding Departments costs per equivalent unit for materials and conRequired 1 Required 2 Required 3 Required 4 Compute the Grinding Departments cost of ending work in process inventory for maRequired 1 Required 2 Required 3 Required 4 Compute the Grinding Departments cost of units transferred out to the Mixing Dep
Answer #1

Unit completed and transferred out = 127,000+352,000-111,000 = 368,000

1) Equivalent unit of production

EUP-Material

EUP-Conversion

Units transferred out

368,000

368,000

Ending work in process

111,000

111,000*70% = 77,700

Equivalent unit of production

479,000

445,700

2) Cost per equivalent unit

Material

Conversion

Beginning work in process

181,610

128,270

Cost added

580,000

375,371

Total Cost

761,610

503,641

Equivalent unit

479,000

445,700

Cost per equivalent unit

1.59

1.13

3) Cost of ending work in process

Material

Conversion

total

Cost of ending work in process

111,000*1.59 = 176,490

77,700*1.13 = 87,801

264,291

4) Cost of units transferred out

Material

Conversion

Total

Cost of units transferred out

368,000*1.59 = 585,120

368,000*1.13 = 415,840

1,000,960

I HOPE IT USEFUL TO YOU IF YOU HAVE ANY DOUBT PLZ COMMENT GIVE ME UP-THUMB. THANKS..